OpenStudy (anonymous):

What is the exact value of tan 195°? −1+3√ / 1+3√ −1+3√ / 1−3√ −1−3√ / 1+3√ −1−3√ / 1−3√

OpenStudy (anonymous):

\[\tan 195=\tan \left( 180+15 \right)=\tan 15=\tan \left( 45-30 \right)=\frac{ \tan 45-\tan 30 }{ 1+\tan 45\tan 30 }\] \[=\frac{ 1-\tan 30 }{ 1+\tan 30 }=?\]

OpenStudy (anonymous):

Define A=180deg. B=15deg. \[\tan (A+B) = \frac{ \tan A + \tan B }{ 1 - \tan A tanB }\] \[\tan A = \tan (180\deg.) = 0\] So \[\tan ( A + B ) = \tan B \] Define B=Θ/2, Θ=30deg. \[\tan (\frac{ \theta }{ 2 }) = \frac{ 1- \cos \theta }{ \sin }\] \[= \frac{ 1- \frac{ \sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } } = 2 - \sqrt{3}\] This problem is not like that, isn't it? \[\frac{ 1 + \sqrt{3} }{ 1 - \sqrt{3} } = \frac{ (1 + \sqrt{3})^{2} }{ 1 - 3 }\] \[\frac{ 4 + \sqrt{3} }{ -2 } = 2 - \sqrt{3}\]